OpenStudy (anonymous):

An unbiased coin is tossed four times. Find the probability of the given event. (Round your answer to four decimal places.) The coin lands heads more than once.

OpenStudy (anonymous):

more than once means two times, three times or four times the probabilities we can find via the fourth line of pascals triangle \[1, 4, 6, 4, 1\] the denominator will be \(2^4=16\) so your probabilities are \[\frac{6}{16},\frac{4}{16}, \frac{1}{16}\] for two, three or four heads respectively add these up